OpenStudy (chaise):

There is a dam with a volume of 49 cubic metres. Water flows out of this into one of the creeks. Salt water with salt content of 30kg/m3 flowed into the dam at 0.68m3/s and mixed with the fresh water. The mixed water flowed out at the same rate. Flag tail in the dam can tolerate salinity up to 14kg/m3. How long after the flow starts can the fish survive in the dam?

OpenStudy (anonymous):

volume = 49 cubic meter salt content = 30 kg/cubic meter rate of water flowed into the dam = 0.68 cubic meter/sec maximum salinity = 14 kg/cubic meter

OpenStudy (anonymous):

wait, solving. just posted the given data to simply my solving :P lol

OpenStudy (anonymous):

oops, "the mixed water flowed out at the same rate". lol

OpenStudy (anonymous):

use derivatives to solve this question. i am weak in derivatives so i quit. lol i

OpenStudy (anonymous):

oh gosh, mechanical engineer!!!

OpenStudy (anonymous):

You can solve for t using V=49m^3, C_0=30kg/m^3, Q=0.68m^3/s and C=14kg/m^3.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I'm not...

OpenStudy (anonymous):

oops,your writing look like those of mechanical engineering students :P

OpenStudy (anonymous):

hehe, messy!